# Context Parametrization with Compositional Adapters

### v2

@@ 4 Experiments
We first outline the experimental design §4.1, with full setup and hyperparameter details in Appendix B. We then evaluate CompAs in two regimes: (i) replacing in-context demonstrations with adapter parameters (§4.2), and (ii) encoding retrieved passages as parametric memory (§4.3). Finally, we assess reconstruction, measuring faithfulness of information encoded in adapter parameters (§4.4). Additional experiments on context order sensitivity and efficiency are in Appendix C.
@@ Methods.
We assess models on end-task performance, measuring accuracy on MMLU and ARC and exact match (EM) and token-level F1 on HotpotQA. Faithfulness is evaluated through the KL divergence between the teacher concatenation and the student sum, as well as token-level F1 (bag-of-tokens). Stability is measured as the standard deviation across ten runs with different sampled contexts. Training follows the weakly supervised protocol from §3.2. Full implementation details, including optimization settings and hyperparameters, are deferred to Appendix B.
@@ 4.2 Encoding Demonstrations as Parameters
We first evaluate CompAs effectiveness in replacing in-context demonstrations with adapter parameters. In the $k$-shot setting ($k\in\{4,8,12,16\}$), demonstrations are partitioned into fixed-size blocks, each block is encoded as an adapter, and the adapters are composed by summing their parameters. Table 1 shows the results on MMLU and ARC-Challenge across three base models.
@@
Llama-3.1 8B Llama-3.2 3B Qwen-2.5 7B Setup / Method MMLU ARC MMLU ARC MMLU ARC 4  ICL $64.2_{1.8}$ $74.2_{1.6}$ $57.2_{1.8}$ $62.9_{1.6}$ $70.2_{1.8}$ $76.5_{1.6}$ 4  GenAda $57.6_{1.8}$ $69.3_{1.7}$ $50.4_{1.4}$ $59.4_{1.7}$ $59.8_{2.1}$ $70.5_{1.7}$ 4  WILDA $\mathbf{68.8}_{0.9}$ $\mathbf{78.2}_{0.7}$ $\mathbf{60.8}_{1.1}$ $\mathbf{66.4}_{1.2}$ $\mathbf{74.0}_{0.4}$ $\mathbf{79.6}_{0.9}$ 4  CompAS $66.7_{0.7}$ $76.5_{0.8}$ $59.0_{0.9}$ $65.1_{0.8}$ $2.4_{0.5} 77$$7.2_{0.6}$ 8  ICL $65.5_{1.7}$ $75.1_{1.5}$ $58.0_{1.7}$ $64.1_{1.5}$ $71.5_{1.7}$ $78.3_{1.5}$ 8  GenAda $59.1_{1.9}$ $69.9_{1.6}$ $51.3_{1.9}$ $61.2_{1.6}$ $60.7_{1.9}$ $71.4_{1.6}$ 8  WILDA $\mathbf{69.8}_{0.7}$ $78.8_{0.8}$ $61.2_{1.1}$ $\mathbf{67.4}_{0.5}$ $\mathbf{74.8}_{0.7}$ $80.0_{0.8}$ 8  CompAS $67.5_{0.5}$ $77.2_{0.9}$ $0.2_{1.3} 66$$6.2_{0.5}$ $73.2_{1.2}$ $78.6_{0.7}$ $2{\times}4$   GenAda $57.8_{1.9}$ $69.5_{1.6}$ $50.9_{0.7}$ $59.4_{1.2}$ $60.5_{1.8}$ $69.2_{1.6}$ $2{\times}4$   WILDA $69.5_{0.3}$ $78.6_{0.9}$ $60.8_{1.1}$ $67.1_{1.0}$ $74.1_{0.4}$ $79.8_{0.8}$ $2{\times}4$CompAS $69.1_{0.7}$ $\mathbf{79.3}_{0.9}$ $\mathbf{61.3}_{1.2}$ $66.8_{0.6}$ $74.6_{0.6}$ $\mathbf{80.2}_{0.9}$ 12  ICL $66.0_{1.7}$ $75.5_{1.5}$ $58.8_{1.7}$ $64.6_{1.5}$ $1.8_{1.7} 77$$7.2_{1.5}$ 12  GenAda $59.9_{1.9}$ $70.2_{1.6}$ $51.6_{1.9}$ $62.5_{1.6}$ $63.1_{1.9}$ $71.6_{1.2}$ 12  WILDA $70.5_{0.8}$ $79.4_{0.6}$ $61.9_{1.1}$ $\mathbf{68.1}_{1.0}$ $75.2_{0.7}$ $79.8_{0.4}$ 12  CompAS $67.2_{1.2}$ $78.9_{0.9}$ $59.5_{1.2}$ $67.2_{0.7}$ $72.9_{1.0}$ $79.3_{1.4}$ $3{\times}4$   GenAda $57.3_{1.9}$ $70.3_{1.6}$ $49.8_{1.9}$ $59.6_{1.6}$ $62.0_{1.9}$ $69.7_{1.6}$ $3{\times}4$   WILDA $70.7_{1.0}$ $79.5_{0.9}$ $62.0_{1.1}$ $67.5_{1.0}$ $75.1_{0.7}$ $79.9_{0.4}$ $3{\times}4$CompAS $\mathbf{71.2}_{1.1}$ $\mathbf{80.1}_{0.5}$ $\mathbf{62.4}_{0.8}$ $67.4_{0.9}$ $\mathbf{75.7}_{0.3}$ $\mathbf{80.3}_{0.8}$ 16  ICL $66.5_{1.7}$ $76.0_{1.5}$ $59.2_{1.7}$ $65.2_{1.5}$ $2.3_{1.7} 77$$7.6_{1.5}$ 16  GenAda $60.5_{1.9}$ $72.5_{1.6}$ $52.7_{1.2}$ $60.8_{1.3}$ $64.2_{1.4}$ $71.9_{0.9}$ 16  WILDA $71.5_{0.7}$ $80.4_{0.3}$ $62.6_{0.5}$ $68.2_{0.3}$ $76.1_{0.4}$ $80.7_{0.8}$ 16  CompAS $68.2_{0.6}$ $79.1_{0.7}$ $62.3_{1.2}$ $68.3_{0.4}$ $73.7_{0.6}$ $79.4_{0.5}$ $4{\times}4$   GenAda $57.6_{1.9}$ $70.6_{1.6}$ $50.3_{1.9}$ $59.9_{1.6}$ $61.3_{1.9}$ $70.4_{1.6}$ $4{\times}4$   WILDA $71.6_{0.6}$ $80.5_{0.9}$ $62.7_{1.1}$ $68.3_{1.1}$ $75.2_{0.5}$ $80.8_{0.8}$ $4{\times}4$CompAS $\mathbf{72.2}_{0.3}$ $\mathbf{81.3}_{0.3}$ $\mathbf{63.4}_{0.7}$ $\mathbf{69.3}_{0.5}$ $\mathbf{77.1}_{0.7}$ $\mathbf{81.5}_{0.4}$
@@
Gains are most pronounced at higher shot counts ($k\in{12,16}$) under $3{\times}4$ and $4{\times}4$ composition settings, where CompAs consistently outperforms wilda (11 out of 12 cases), despite the latter’s use of context-specific adapter fine-tuning. We hypothesize that this advantage arises from an implicit chunking–composition mechanism, where the model partitions demonstrations into manageable subsets. Each adapter encodes a partial context view, and their addition reconstructs the effect of all demonstrations. As a result, CompAs scales gracefully with the number of demonstrations: it leverages composition to maintain strong performance in long contexts, ultimately matching or exceeding specialized adapter fine-tuning while retaining efficiency and composability.
@@ 4.3 Encoding Context as Parametric Memory
CompAs consistently outperforms the teacher (base LLM) across all settings. Absolute performance is lower on pairs and triplets – since the model must retain context for multiple queries while being evaluated on each one separately – but the relative gains over the teacher are larger in these harder settings. This indicates that CompAs is particularly effective at composing and retaining information across multiple contexts.
@@
Setup Method LLaMA-3.1 8B LLaMA-3.2 3B Qwen-2.5 7B Single ($\mathbf{c}$) Teacher $70.9/82.0$ $65.9/78.1$ $69.2/80.7$ CompAS $72.2/82.9$ $67.0/78.6$ $69.4/82.2$ Pair ($\mathbf{c}_{1};\mathbf{c}_{2}]$) Teacher (concat) $69.0/80.8$ $64.1/76.8$ $67.5/79.8$ CompAS (concat) $69.4/81.1$ $64.5/77.0$ $67.9/79.5$ CompAS (composed) $71.3/82.3$ $66.2/78.4$ $69.6/80.9$ Triplet ($\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{c}_{3}]$) Teacher (concat) $66.3/78.5$ $61.2/74.1$ $64.0/76.0$ CompAS (concat) $67.5/79.4$ $62.0/75.3$ $65.1/77.0$ CompAS (composed) $71.1/82.0$ $64.8/76.9$ $68.1/80.4$
@@ 4.4 Context Reconstruction
Finally, we test whether composed adapters preserve the informational content of their supports by prompting the model to reconstruct missing contexts. Using a special [RECON] token, we decode from adapters and compute token-level F1 on MMLU, ARC, and HotpotQA. Table 3 shows that CompAs maintains consistently high reconstruction fidelity, with only small drops as the context expands to pair and triplet units.
@@
Llama-3.1 8B Llama-3.2 3B Qwen-2.5 7B Setup MMLU ARC Hotpot MMLU ARC Hotpot MMLU ARC Hotpot Single (1 unit) $91.2$ $89.5$ $87.8$ $89.0$ $87.2$ $85.4$ $90.1$ $88.1$ $86.2$ Pair (2 units) $90.3$ $88.9$ $86.0$ $88.1$ $86.3$ $83.5$ $89.4$ $87.5$ $84.3$ Triplet (3 units) $88.4$ $86.9$ $83.7$ $86.5$ $84.8$ $81.9$ $87.6$ $85.9$ $82.5$
@@ 5 Analysis
Building on the effectiveness of CompAs, we analyze the role of generator capacity (§5.1) and the contribution of each loss component (§5.2). We discuss limitations of our approach in Appendix E.
@@ 5.1 Capacity and Weak-to-Strong Generalization
We explore three generator capacities $G$: Adapter (CompAs) – corresponds to our default setup (§3); RNN – a lightweight recurrent network which aggregates token-level encodings of $\mathbf{c}$ before the same linear head produces $G(\mathbf{c})$; Linear – an ablation that removes the generator adapter and recurrent aggregator, leaving a single bottleneck linear projection of a pooled context representation (see configuration details in Appendix B).
@@
In Figure 3, we analyze the effects of W2S generalization by grouping MMLU tasks into five difficulty levels (1155). To obtain these groups, we evaluate the base LLaMA 8B with $16$ shots on each MMLU subtask and sort them by accuracy, partitioning into five bins. We then evaluate generators in a $4{\times}4$ setup ($16$ demonstrations in total). Across levels 1133, all generator variants improve over ICL, with gains ordered by capacity (Adapter, RNN, and Linear). At higher difficulty, differences sharpen: the linear generator collapses at level 4455, RNN stagnates and drops at level $5$, while Adapter continues to improve, achieving the largest gains on the hardest tasks. These results show that sufficient generator capacity is crucial for surpassing the ICL teacher and maintaining additive composition as complexity grows. CompAs scales with task difficulty, enhancing W2S generalization on the hardest tasks.
@@
Figure 3: Accuracy deltas of different generators over ICL across five MMLU difficulty buckets (average over $10$ runs for LLaMA 8B).
@@ 5.2 Importance of Loss Components
We always include $\mathcal{L}_{\text{ST}}$, since weak supervision from the teacher is essential for transferring contextual information into the student. Dropping $\mathcal{L}_{\text{ADD}}$ or $\mathcal{L}_{\text{RECON}}$ degrades performance and shows that they play complementary roles. $\mathcal{L}_{\text{ADD}}$ reduces additivity error ($\eta$) and stabilizes multi-context composition. $\mathcal{L}_{\text{RECON}}$ provides smaller but consistent gains while also enabling context reconstruction. Removing both terms ($\mathcal{L}_{\text{ADD}}$ and $\mathcal{L}_{\text{RECON}}$) leads to the most pronounced drop, confirming that they jointly support the effectiveness of the student. Overall, the full objective yields the strongest results, consistent with the theory in §2.
Table 4: Ablation of loss function components. Results are on MMLU and ARC ($4{\times}4$ demos) and HotpotQA (EM/F1, pairs). We report averages over $10$ seeds with standard deviations as subscripts.
Objective Variant MMLU ARC-Challenge HotpotQA (EM/F1)

### v3

@@ 4 Experiments
We first outline the experimental design Section 4.1, with full setup and hyperparameter details in Appendix B. We then evaluate CompAs in two regimes: (i) replacing in-context demonstrations with adapter parameters (Section 4.2), and (ii) encoding retrieved passages as parametric memory (Section 4.3). Finally, we assess reconstruction, measuring faithfulness of information encoded in adapter parameters (Section 4.4). Additional experiments on context order sensitivity and efficiency are in Appendix C.
@@ Methods.
We assess models on end-task performance, measuring accuracy on MMLU and ARC and exact match (EM) and token-level F1 on HotpotQA. Faithfulness is evaluated through the KL divergence between the teacher concatenation and the student sum, as well as token-level F1 (bag-of-tokens). Stability is measured as the standard deviation across ten runs with different sampled contexts. Training follows the weakly supervised protocol from Section 3.2. Full implementation details, including optimization settings and hyperparameters, are deferred to Appendix B.
@@ 4.2 Encoding Demonstrations as Parameters
We first evaluate effectiveness in replacing in-context demonstrations with adapter parameters. In the $k$-shot setting ($k\in\{4,8,12,16\}$), demonstrations are partitioned into fixed-size blocks, each block is encoded as an adapter, and the adapters are composed by summing their parameters. Table 1 shows the results on MMLU and ARC-Challenge across three base models.
@@
Llama-3.1 8B Llama-3.2 3B Qwen-2.5 7B
Setup / Method MMLU ARC MMLU ARC MMLU ARC
4    ICL $64.2_{1.8}$ $74.2_{1.6}$ $57.2_{1.8}$ $62.9_{1.6}$ $70.2_{1.8}$ $76.5_{1.6}$
4    GenAda $57.6_{1.8}$ $69.3_{1.7}$ $50.4_{1.4}$ $59.4_{1.7}$ $59.8_{2.1}$ $70.5_{1.7}$
4    wilda $\mathbf{68.8}_{0.9}$ $\mathbf{78.2}_{0.7}$ $\mathbf{60.8}_{1.1}$ $\mathbf{66.4}_{1.2}$ $\mathbf{74.0}_{0.4}$ $\mathbf{79.6}_{0.9}$
4    CompAs $66.7_{0.7}$ $76.5_{0.8}$ $59.0_{0.9}$ $65.1_{0.8}$ $2.4_{0.5} 77$$7.2_{0.6}$
8    ICL $65.5_{1.7}$ $75.1_{1.5}$ $58.0_{1.7}$ $64.1_{1.5}$ $71.5_{1.7}$ $78.3_{1.5}$
8    GenAda $59.1_{1.9}$ $69.9_{1.6}$ $51.3_{1.9}$ $61.2_{1.6}$ $60.7_{1.9}$ $71.4_{1.6}$
8    wilda $\mathbf{69.8}_{0.7}$ $78.8_{0.8}$ $61.2_{1.1}$ $\mathbf{67.4}_{0.5}$ $\mathbf{74.8}_{0.7}$ $80.0_{0.8}$
8    CompAs $67.5_{0.5}$ $77.2_{0.9}$ $0.2_{1.3} 66$$6.2_{0.5}$ $73.2_{1.2}$ $78.6_{0.7}$
$2{\times}4$   GenAda $57.8_{1.9}$ $69.5_{1.6}$ $50.9_{0.7}$ $59.4_{1.2}$ $60.5_{1.8}$ $69.2_{1.6}$
$2{\times}4$wilda $69.5_{0.3}$ $78.6_{0.9}$ $60.8_{1.1}$ $67.1_{1.0}$ $74.1_{0.4}$ $79.8_{0.8}$
$2{\times}4$CompAS $69.1_{0.7}$ $\mathbf{79.3}_{0.9}$ $\mathbf{61.3}_{1.2}$ $66.8_{0.6}$ $74.6_{0.6}$ $\mathbf{80.2}_{0.9}$
12    ICL $66.0_{1.7}$ $75.5_{1.5}$ $58.8_{1.7}$ $64.6_{1.5}$ $1.8_{1.7} 77$$7.2_{1.5}$
12    GenAda $59.9_{1.9}$ $70.2_{1.6}$ $51.6_{1.9}$ $62.5_{1.6}$ $63.1_{1.9}$ $71.6_{1.2}$
12    wilda $70.5_{0.8}$ $79.4_{0.6}$ $61.9_{1.1}$ $\mathbf{68.1}_{1.0}$ $75.2_{0.7}$ $79.8_{0.4}$
12    CompAs $67.2_{1.2}$ $78.9_{0.9}$ $59.5_{1.2}$ $67.2_{0.7}$ $72.9_{1.0}$ $79.3_{1.4}$
$3{\times}4$   GenAda $57.3_{1.9}$ $70.3_{1.6}$ $49.8_{1.9}$ $59.6_{1.6}$ $62.0_{1.9}$ $69.7_{1.6}$
$3{\times}4$wilda $70.7_{1.0}$ $79.5_{0.9}$ $62.0_{1.1}$ $67.5_{1.0}$ $75.1_{0.7}$ $79.9_{0.4}$
$3{\times}4$CompAs $\mathbf{71.2}_{1.1}$ $\mathbf{80.1}_{0.5}$ $\mathbf{62.4}_{0.8}$ $67.4_{0.9}$ $\mathbf{75.7}_{0.3}$ $\mathbf{80.3}_{0.8}$
16    ICL $66.5_{1.7}$ $76.0_{1.5}$ $59.2_{1.7}$ $65.2_{1.5}$ $2.3_{1.7} 77$$7.6_{1.5}$
16    GenAda $60.5_{1.9}$ $72.5_{1.6}$ $52.7_{1.2}$ $60.8_{1.3}$ $64.2_{1.4}$ $71.9_{0.9}$
16    wilda $71.5_{0.7}$ $80.4_{0.3}$ $62.6_{0.5}$ $68.2_{0.3}$ $76.1_{0.4}$ $80.7_{0.8}$
16    CompAs $68.2_{0.6}$ $79.1_{0.7}$ $62.3_{1.2}$ $68.3_{0.4}$ $73.7_{0.6}$ $79.4_{0.5}$
$4{\times}4$   GenAda $57.6_{1.9}$ $70.6_{1.6}$ $50.3_{1.9}$ $59.9_{1.6}$ $61.3_{1.9}$ $70.4_{1.6}$
$4{\times}4$wilda $71.6_{0.6}$ $80.5_{0.9}$ $62.7_{1.1}$ $68.3_{1.1}$ $75.2_{0.5}$ $80.8_{0.8}$
$4{\times}4$CompAs $\mathbf{72.2}_{0.3}$ $\mathbf{81.3}_{0.3}$ $\mathbf{63.4}_{0.7}$ $\mathbf{69.3}_{0.5}$ $\mathbf{77.1}_{0.7}$ $\mathbf{81.5}_{0.4}$
@@
Gains are most pronounced at higher shot counts ($k\in{12,16}$) under $3{\times}4$ and $4{\times}4$ composition settings, where CompAs consistently outperforms wilda (11 out of 12 cases), despite the latter’s use of context-specific adapter fine-tuning. We hypothesize that this advantage arises from an implicit chunking–composition mechanism, where the model partitions demonstrations into manageable subsets. Each adapter encodes a partial context view, and their addition reconstructs the effect of all demonstrations. As a result, CompAs scales gracefully with the number of demonstrations: it leverages composition to maintain strong performance in long contexts, ultimately matching or exceeding specialized adapter fine-tuning while retaining efficiency and composability. A more detailed analysis of how block granularity affects compositional stability and scaling behavior is provided in Appendix C.4.
@@ 4.3 Encoding Context as Parametric Memory
CompAs consistently outperforms the teacher across all settings. Absolute performance is lower on pairs and triplets – since the model must retain context for multiple queries while being evaluated on each one separately – but the relative gains over the teacher are larger in these harder settings. This indicates that CompAs is particularly effective at composing and retaining information across multiple contexts.
@@
Setup Method LLaMA-3.1 8B LLaMA-3.2 3B Qwen-2.5 7B
Single ($\mathbf{c}$) Teacher $70.9/82.0$ $65.9/78.1$ $69.2/80.7$
CompAS $72.2/82.9$ $67.0/78.6$ $69.4/82.2$
Pair ($\mathbf{c}_{1};\mathbf{c}_{2}]$) Teacher (concat) $69.0/80.8$ $64.1/76.8$ $67.5/79.8$
CompAS (concat) $69.4/81.1$ $64.5/77.0$ $67.9/79.5$
CompAS (composed) $71.3/82.3$ $66.2/78.4$ $69.6/80.9$
Triplet ($\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{c}_{3}]$) Teacher (concat) $66.3/78.5$ $61.2/74.1$ $64.0/76.0$
CompAS (concat) $67.5/79.4$ $62.0/75.3$ $65.1/77.0$
CompAS (composed) $71.1/82.0$ $64.8/76.9$ $68.1/80.4$
@@ 4.4 Context Reconstruction
Finally, we test whether composed adapters preserve the informational content of their supports by prompting the model to reconstruct missing contexts. Using a special [RECON] token, we decode from adapters and compute token-level F1 on MMLU, ARC, and HotpotQA. Table 3 shows that CompAs maintains consistently high reconstruction fidelity, with only small drops as the context expands to pair and triplet units. In Section 5.2, we further show that ablating the reconstruction objective leads to a consistent drop in downstream performance, highlighting its importance beyond faithfulness alone.
@@
Llama-3.1 8B Llama-3.2 3B Qwen-2.5 7B
Setup MMLU ARC Hotpot MMLU ARC Hotpot MMLU ARC Hotpot
Single (1 unit) $91.2$ $89.5$ $87.8$ $89.0$ $87.2$ $85.4$ $90.1$ $88.1$ $86.2$
Pair (2 units) $90.3$ $88.9$ $86.0$ $88.1$ $86.3$ $83.5$ $89.4$ $87.5$ $84.3$
Triplet (3 units) $88.4$ $86.9$ $83.7$ $86.5$ $84.8$ $81.9$ $87.6$ $85.9$ $82.5$
Beyond accuracy, compositionality, and reconstruction fidelity, we also examine the computational footprint of CompAs. A detailed efficiency analysis is provided in Appendix C.6, where we compare inference FLOPs, peak memory, and cumulative training cost across methods. The results show that CompAs decouples inference cost from context length, yielding increasing speedups at higher shot counts while maintaining a moderate amortized training cost.
@@ 5 Analysis
Building on the effectiveness of CompAs, we analyze the role of generator capacity (Section 5.1) and the contribution of each loss component (Section 5.2). A broader discussion of the limitations of our approach is deferred to Appendix E.
@@ 5.1 Capacity and Weak-to-Strong Generalization
We explore three generator capacities $G$: Adapter (CompAs) – corresponds to our default setup (Section 3); RNN – a lightweight recurrent network which aggregates token-level encodings of $\mathbf{c}$ before the same linear head produces $G(\mathbf{c})$; Linear – an ablation that removes the generator adapter and recurrent aggregator, leaving a single bottleneck linear projection of a pooled context representation (see configuration details in Appendix B).
@@
In Figure 3, we analyze the effects of W2S generalization by grouping MMLU tasks into five difficulty levels (1155). To obtain these groups, we evaluate the base LLaMA 8B with $16$ shots on each MMLU subtask and sort them by accuracy, partitioning into five bins. We then evaluate generators in a $4{\times}4$ setup ($16$ demonstrations in total). Across levels 1133, all generator variants improve over ICL, with gains ordered by capacity (Adapter, RNN, and Linear). At greater difficulty, differences sharpen: the linear generator collapses at level 4455, RNN stagnates and drops at level $5$, while Adapter continues to improve, achieving the largest gains on the hardest tasks. These results show that sufficient generator capacity is crucial for surpassing the ICL teacher and maintaining additive composition as complexity grows. CompAs scales with task difficulty, enhancing W2S generalization on the hardest tasks.
@@
Figure 3: Accuracy deltas of different generators relative to ICL across five MMLU difficulty buckets (stratified by base LLM). Results are averaged over $10$ runs using LLaMA 8B.
@@ 5.2 Importance of Loss Components
We always include $\mathcal{L}_{\text{ST}}$, since weak supervision from the teacher is essential for transferring contextual information into the student. Dropping $\mathcal{L}_{\text{ADD}}$ or $\mathcal{L}_{\text{RECON}}$ degrades performance and shows that they play complementary roles. $\mathcal{L}_{\text{ADD}}$ reduces additivity error ($\eta$) and stabilizes multi-context composition. $\mathcal{L}_{\text{RECON}}$ provides smaller but consistent gains while also enabling context reconstruction. Removing both terms ($\mathcal{L}_{\text{ADD}}$ and $\mathcal{L}_{\text{RECON}}$) leads to the most pronounced drop, confirming that they jointly support the effectiveness of the student. Overall, the full objective yields the strongest results, consistent with the theory in Section 2.
Table 4: Ablation of loss function components. Results are on MMLU and ARC ($4{\times}4$ demos) and HotpotQA (EM/F1, pairs). We report averages over $10$ seeds.
Objective MMLU ARC HotpotQA (EM/F1)